having trouble with this problem,
(2t-5)(t-1)=2
my book calls this part of the chapter "solving equations by factoring", if that helps you any.

Answer by stanbon(48579)

(2t-5)(t-1)=2
2t^2-7t+5=2
2t^2-7t+3=0
(2t-1)(t-3)=0
t=1/2 or t=3
Cheers,
Stan H.